### v2

Abstract
Gravitational particle production in cosmology is a mechanism through which particles of different natures are produced during the very early universe. It is a general mechanism that explains how the universe became populated with the particles of the Standard Model after cosmic inflation and may also account for the origin of dark matter. In this work, we study the non-perturbative production of massive scalar particles in the Higgs-$R^{2}$ inflation model, a two-field scalar inflation model within the Einstein frame. We consider spectator scalar fields that are conformally or minimally coupled to the gravitational field through the curvature scalar $R$ which in turn is a time-dependent function determined by the fields driving inflationary dynamics. We numerically compute the production of these particles using the Bogolyubov transformation method for each scenario, aiming to assess the spectrum of the produced particles. For both scenarios, we consider light particles with masses $m_{\chi}\ll H_{\mathrm{end}}$ and large masses that exceed the Hubble scale at the end of inflation $m_{\chi}\gtrsim H_{\mathrm{end}}$.
keywords:
Particle production, Multi-field inflation, Higgs inflation, Early universe
1 Introduction
Cosmic inflation is the cornerstone of modern cosmology, providing a framework to understand the early universe and predict its fundamental characteristics.
It describes a period of accelerated expansion driven by a scalar field $\phi$ that slowly rolls towards the minimum of its potential $V(\phi)$. This mechanism addresses classical problems of the standard cosmological model and has become a central paradigm in cosmology [1, 2]. During the inflationary epoch, significant particle production occurs, particularly when the kinetic energy of the inflaton field becomes comparable to its potential energy. This particle production can arise from gravitational effects due to the universe’s expansion and the dynamics of the inflaton field, serving as a classical background for propagating quantum fields [3, 4]. This phenomenon, known as gravitational particle production, is a fundamental mechanism in quantum field theory in curved spacetime, offering insights into the origins of dark matter and the preheating and reheating epochs following inflation [5]. Several scenarios have been proposed for vacuum particle production during the early universe, particularly during inflation due to the evolution of the inflaton field and the universe’s expansion. Typically, the inflaton field is coupled to matter fields (bosonic, fermionic, or gauge fields) by a coupling constant $g$. At a critical value $\phi=\phi_{*}$ of the inflaton field, resonance is triggered, leading to instantaneous particle production [6, 7]. However, matter fields need not be directly coupled with the inflaton or standard model particles. Massive particles can be produced purely through gravitational effects, either via direct couplings with the gravitational field or induced couplings through a conformal transformation. If the contribution of the matter field to the energy density is negligible, the inflationary dynamics remain unaffected, leading to no back-reaction effects on spacetime geometry. Such fields are called spectator fields, and this mechanism provides a natural explanation for the origin of dark matter [8, 9, 10, 11, 12, 13]. Beyond its role in the origin of dark matter, gravitational particle production is also linked to key early-universe phenomena such as preheating after inflation [14, 15, 16, 17], primordial gravitational waves [18, 19, 20], and primordial magnetogenesis [21, 22, 23, 24]. Furthermore, it leaves significant observable imprints on the Cosmic Microwave Background (CMB), particularly in the generation of primordial non-Gaussianities [25, 26, 27].
In this work, we study gravitational particle production during and immediately after inflation driven by the SM Higgs boson and the scalaron field in the Higgs-$R^{2}$ inflation model. This inflation model extends the SM Higgs inflationary model [28] to address the problem of large coupling [29, 30]. Here, cosmic inflation is driven by the Higgs field and the extra scalar degree of freedom from the quadratic $R$ term, known as the scalaron. Unlike other multi-field inflation models, the Higgs-$R^{2}$ model effectively reduces to a single-field model during inflation, with the Higgs field acting as an auxiliary field while the scalaron drives inflation [31]. At the end of inflation during the reheating phase, both fields and the Ricci scalar oscillate around the potential minimum. This oscillatory behavior enhances scalar particle production through non-perturbative mechanisms such as tachyonic instabilities [32, 33]. We propose an extension to the Higgs-$R^{2}$ model by introducing a massive scalar field $\chi$ as a spectator field that does not contribute to the energy density during the inflationary epoch, with negligible interactions with the SM but considering a direct coupling to the Ricci scalar via a non-minimal coupling term $\xi_{\chi}R\,\chi^{2}$. Initially, the inflationary sector and the spectator field are not explicitly coupled in the Jordan frame; however, gravitational interactions between the scalaron, the Higgs field, and free scalars induce couplings upon transforming to the Einstein frame. This process results in gravitational particle production with significant implications for preheating after inflation and background dynamics. Induced coupling in the Einstein frame between free scalars and the scalar sector of the action provides an excellent channel for scalaron decay into light scalars. However, in the non-perturbative approach, the field mass $m_{\chi}$ does not necessarily have to be light, and we can explore the possibility of scalar particle production for spectator field masses $m_{\chi}\gtrsim H_{\mathrm{end}}$, where $H_{\mathrm{end}}$ is the Hubble ratio at the end of inflation. We take into account the oscillatory behavior of background dynamics after inflation as a classical background, which results in particle production due to the time-dependent background inducing a time-dependent frequency in the mode equations for the spectator field $\chi$. To achieve this, we employ the Bogoliubov approach [3] to compute the comoving number density $n_{k}(t)$ of particles generated by the background dynamics in Higgs-$R^{2}$ inflation, solving the mode equations for the quantized modes $\chi$.
The article is organized as follows: Section 2 provides a brief introduction to the Higgs-$R^{2}$ model and the background dynamics during and after inflation for a set of suitable initial conditions. Section 3 analyzes the couplings induced by gravitational effects in the Einstein frame, in Section 4, we derive the mode equation for the spectator field, which is necessary to investigate the behavior of the time-dependent frequency during and after inflation. We adopt a numerical approach to solve the mode equation for modes that leave 5 e-folds before the horizon crossing during inflation, ensuring that each mode is within the horizon at the start of the numerical simulation. Section 4 examines the production of massive particles during preheating. Finally, Section 5 discusses the results and conclusions.
In the following, we will use natural units in which $\hbar=1$, $c=1$, $\kappa_{\mathrm{B}}=1$ and we use the mass Planck as $8\pi G=M_{p}^{-2}$. We also work in the context of a spatially flat, homogeneous, and isotropic universe described by the FLRW metric
$\mathrm{d}s^{2}$ $=$ $-\mathrm{d}t^{2}+a^{2}(t)(\mathrm{d}x^{2}+\mathrm{d}y^{2}+\mathrm{d}z^{2})$
$=$ $a^{2}(\tau)(-\mathrm{d}\tau^{2}+\mathrm{d}x^{2}+\mathrm{d}y^{2}+\mathrm{d}z^{2})\,,$
where $t$ is the cosmic time, $\tau$ the conformal time and the relation between the two is $\mathrm{d}\tau=\mathrm{d}t/a(t)$ with $a(t)$ is the scale factor.
2 Brief review of Higgs-$R^{2}$ inflationary model
In this section, we will give a brief review of the Higgs-$R^{2}$ model in both Einstein and Jordan frames. For further details, we recommend the main references [31, 34, 30, 29] and the references therein. Let us get started with the discussion in the Jordan frame; the model is given by the action
$S_{\mathrm{J}}=\int\mathrm{d}x^{4}\sqrt{-g_{J}}\left[\dfrac{1}{2}(M_{p}^{2}+\xi_{h}h^{2})R_{J}+\dfrac{\xi_{s}}{4}\,R_{J}^{2}-\dfrac{1}{2}g^{\mu\nu}_{J}\,\partial_{\mu}h\,\partial_{\nu}h-\dfrac{\lambda}{4}\,h^{4}\right]\,,$ (2)
where $h$ is the Higgs field in the unitary gauge, the subindex $J$ denotes variables defined in the Jordan frame, $\xi_{h}$ is the constant coupling between the Higgs field and the gravitational sector, $\xi_{s}$ is a mass parameter and $\lambda$ is the quartic self constant coupling of the Higgs field. This model contains a quadratic term in $R_{J}$ described by the constant $\xi_{s}$, which comes from the $R^{2}$ model [35]. Such a constant is related to the scalaron mass as $\xi_{s}=M_{p}^{2}/3M^{2}$. The model (2) can be regarded as a UV extension of Higgs inflation, so it is interesting to make a double interpretation of the gravitational coupling in this conformal frame. In Higgs inflation, the constant coupling $\xi_{h}$ must fulfill the condition $\xi_{h}^{2}/\lambda\sim 10^{10}$ to be consistent with CMB observations [28, 36]. This causes the cutoff scale of the model to be below the natural energy scale (Planck scale) during inflation, making the predictions of the model questionable; however, it has been shown that including the Ricci scalar quadratic term modifies the value of the energy scale, raising it to the Planck scale. To solve the strong coupling problem present in the scalar sector, the parameter $\xi_{s}$ has to obey the relation [29]
$\xi_{s}\gtrsim\dfrac{\xi_{h}^{2}}{4\pi}\,.$ (3)
This expression allows us to make a perturbative treatment of the model. To describe the inflationary dynamics of the model, it is more convenient to bring the action (2) into the canonical Einstein-Hilbert form. So we need to carry out a conformal transformation $\Omega^{2}(x)$ introducing an extra scalar degree of freedom $\phi$ called scalaron, which comes from the quadratic term in $R_{J}$. Thus, to go from the Jordan frame to Einstein frame one has to perform a conformal transformation of the space-time metric $g_{\mu\nu}\to\Psi^{2}(x)g_{\mu\nu}$ with $\Psi(x)>0$, where the correct form of it is
$\phi=\alpha^{-1}\log\Psi^{2}\,,$ (4)
where $\alpha=\sqrt{2/3}M_{p}^{-1}$. The conformal transformation (4) removes the quadratic term of the scalar curvature $R_{J}$ introducing the scalaron $\phi$. Then in the Einstein frame, the Higgs-$R^{2}$ model of inflation takes the form of a two-scalar fields model with a non-canonical kinetic term
$S_{E}=\int\mathrm{d}^{4}x\,\sqrt{-g}\,\left\{\dfrac{M_{p}^{2}}{2}R-\dfrac{1}{2}g^{\mu\nu}\partial_{\mu}\phi\,\partial_{\nu}\phi-\dfrac{1}{2}\,e^{-\alpha\phi}g^{\mu\nu}\partial_{h}\,\partial_{\nu}h-V(\phi\,,h)\right\}\,,$ (5)
which can be written in the general action for multi-field inflationary models
Figure 1: Scalar potential in the Einstein frame of the Higgs-$R^{2}$ model seen from different angles. The curve $\phi=\mathbf{cte}$ has the shape of the potential of the $R^{2}$ inflation model, while $h=\mathrm{cte}$ has the shape of the quartic Higgs potential of the Standard Model. We have restricted ourselves to the standard parameter space $\xi_{h}=4000$, $\xi_{s}=2\times 10^{9}$ and $\lambda=0.013$.
$S=\int\mathrm{d}^{4}x\,\sqrt{-g}\left[\dfrac{M_{p}^{2}}{2}R-\dfrac{1}{2}G_{IJ}\,g^{\mu\nu}\partial_{\mu}\phi^{I}\,\partial_{\nu}\phi^{J}-V(\phi^{I})\right]\,,$ (6)
where in our case $\phi^{I}=(\phi\,,h)$ and $G_{IJ}$ is the non-flat field-space metric given by
$G_{IJ}=\begin{pmatrix}1&0\\ 0&e^{-\alpha\phi}\end{pmatrix}\,.$ (7)
We can see that the non-canonical kinetic term is associated with the geometry of the field space via the metric $G_{IJ}$ [37]. On the other hand, the potential in the Einstein frame has interaction terms between the Higgs field $h$ and the scalaron $\phi$ and takes the form
$V(\phi\,,h)=\dfrac{1}{4}\,e^{-2\alpha\phi}\left[\dfrac{M_{p}^{4}}{\xi_{s}}\left(e^{\alpha\phi}-1-\dfrac{\xi_{h}}{M_{p}^{2}}\,h^{2}\right)^{2}+\lambda\,h^{4}\right]\,.$
The potential $V(\phi^{I})$ depends on both the $\phi$ and $h$ and mixes terms which make it more difficult to analyze analytically the background trajectory. Nevertheless, during inflation, we can make an effective approximation of the model such that inflation is driven by the single scalar field $\phi$ [31] and with the following interesting features
  1. 1.
    $V(\phi^{I})$ shows two valleys and one ridge.
  2. 2.
    For $\xi_{h}$, $\xi_{s}$ positive and enough large, the potential presents an attractor behavior.
  3. 3.
    If $\phi\gg M_{p}$ with arbitrary initial conditions, the fields $\phi$ and $h$ will always evolve towards one of the valleys thus beginning the slow-roll evolution.
  4. 4.
    If $\phi\ll M_{p}$, the potential is reduced to Higgs potential
    $V(\phi\ll M_{p}\,,h)\simeq V(h)=\dfrac{\lambda_{\mathrm{eff}}}{4}\,h^{4}\,,$ (9)
    where $\lambda_{\mathrm{eff}}$ is the effective quartic self-coupling constant defined by
    $\lambda_{\mathrm{eff}}=\lambda+\dfrac{\xi_{h}^{2}}{\xi_{s}}\,.$ (10)
    The condition $\lambda_{\mathrm{eff}}\lesssim 4\pi$ is necessary for perturbative preheating
  5. 5.
    If $h\ll M_{p}$, the potential is reduced to Starobinsky potential
    $V(\phi\,,h\ll M_{p})\simeq\dfrac{M_{p}^{4}}{4\xi_{s}}(1-e^{-\alpha\phi})^{2}\,.$ (11)
We can see the positions of the valleys and the hill at $h=0$ in figure 1. The inflationary dynamics occur along one of these valleys, showing a universal attractor behavior, so that inevitably the fields $\phi$ and $h$ will always fall into one of the valleys and then drive slow-roll inflation. To find the valleys, we can perform the analysis in both Einstein [31] and Jordan [38] frames. In the Einstein frame in the limit $\phi\gg M_{p}$ (it is required for enough long inflation) and $h\gg v$, where $v=246\,\mathrm{GeV}$, the valleys are found out by the condition
$\dfrac{\partial V}{\partial h}=0\,,$ (12)
which implies that the Higgs field is completely determined by the inflaton field $\phi$
$h^{2}_{\mathrm{min}}(\phi)=\dfrac{M_{p}^{2}\xi_{h}}{\xi_{h}^{2}+\lambda\xi_{s}}\,(e^{\alpha\phi}-1)\,.$ (13)
Because the potential (8) is symmetric with respect to transformation $h\to-h$, the two valleys are localized by the condition (13). Thus, this shows that the Higgs field is an auxiliary field during inflation, so if we put the result (13) back into the Einstein potential (8), we give the effective potential along the valleys
$V_{\mathrm{eff}}(\phi\,,h_{\mathrm{min}})=\dfrac{M_{p}^{4}}{4(\xi_{s}+\xi_{h}^{2}/\lambda)}\,(1-e^{-\alpha\phi})^{2}$ (14)
Notice that the effective potential reduces to Higgs inflation if $\lambda\xi_{s}\ll\xi_{h}^{2}$ while one gets $R^{2}$ inflation for $\lambda\xi_{s}\gg\xi_{h}^{2}$. To provide the correct amplitude of primordial scalar perturbations, the amplitude of (14) has to fulfill [39]
$\xi_{s}+\dfrac{\xi_{h}^{2}}{\lambda}=\dfrac{N_{*}^{2}}{72\pi^{2}A_{\zeta}}\simeq 2\times 10^{9}\,,$ (15)
where $N_{*}\sim 50-60$ is the number of e-folds before the end of inflation, and $A_{\zeta}\sim 2\times 10^{-9}$ is the amplitude of the primordial power spectrum. For $\lambda=0.013$ and $\xi_{h}\,,\xi_{s}>0$ which satisfy (15), the inflationary dynamics is effectively the same as single-field slow-roll inflation. In this paper, we will only analyze the standard parameter region $\xi_{h}\,,\xi_{s}\,,\lambda>0$. Specifically, if we have $\xi_{h}\gg 0$ the isocurvature effects of the model are absent the bending is within the weak mixing case $\eta_{\perp}\to 0$ [31], and the isocurvature and adiabatic modes are not coupled so that they evolve independently in the standard parameter region of Higgs-$R^{2}$ model. Therefore, the description of the dynamics inside the valleys is an effective single-field inflation.
2.1 Background dynamics
What follows is to analyze the dynamics of the model. We assume that during inflation both scalaron $\phi$ and Higgs field $h$ are homogeneous and only depend on cosmic time $t$. In this case, the equations of motion can be written as
$\ddot{\phi}+3H\dot{\phi}+\dfrac{\partial V}{\partial\phi}$ $=$ $-\dfrac{\alpha}{2}\,e^{-\alpha\phi}\dot{h}^{2}\,,$ (16)
$\ddot{h}+3H\dot{h}+e^{\alpha\phi}\,\dfrac{\partial V}{\partial h}$ $=$ $\alpha\dot{\phi}\,\dot{h}\,.$ (17)
The Einstein equations give the Friedman equations for $H(t)=\dot{a}/a$
$3M_{p}^{2}\,H^{2}$ $=$ $\dfrac{1}{2}\dot{\phi}^{2}+\dfrac{1}{2}\,e^{-\alpha\phi}\dot{h}^{2}+V(\phi\,,h)\,,$ (18)
$\dot{H}$ $=$ $-\dfrac{1}{2M_{p}^{2}}\,\left(\dot{\phi}^{2}+e^{-\alpha\phi}\dot{h}^{2}\right)\,.$ (19)
Figure 2: We present the numerical solution of the background dynamics after inflation in cosmic time $t$ for the Higgs and scalaron fields within the valley of the potential. We have numerically solved (16), (17), and (18) using the numerical parameters $\xi_{h}=4000$, $\xi_{s}=2\times 10^{9}$ and $\lambda=0.013$ together with initial conditions $\phi(t_{0})=5.7M_{p}$ and $h(t_{0})=0.096M_{p}$.
One has to numerically solve the system of equations (16), (17), (18) and (19) to explore the inflationary dynamics of the model in the Einstein frame given some combination values of the parameters $\xi_{h}\,,\xi_{s}\,,\lambda$ and given suitable initial conditions for $\phi(t_{0})$, $h(t_{0})$. The inflation occurs when $\phi\gg M_{p}$ and the value of the auxiliary field $h$ is given by (13) inside one of the valleys of the potential. In the figure 2 we present the oscillatory regime of both fields after inflation in terms of cosmic time for the initial conditions $\phi_{0}=5.7M_{p}$ and $h_{0}\approx 0.096M_{p}$. However, according to the argument we have presented above, during inflation the isocurvature effects are absent and the dynamics inside the valleys are an effective single-field description. Then the scalaron field should fulfill the single-field equation
$\ddot{\phi}+3H\dot{\phi}+\dfrac{\mathrm{d}V_{\mathrm{eff}}}{\mathrm{d}\phi}\simeq 0\,,$ (20)
and the evolution of the Higgs field is determined by the condition (13). We can assume a single-field slow roll approximation in the inflationary dynamics inside the valleys. This happens when $\ddot{\phi}\ll 3H\dot{\phi}$, thus the field equation for $\phi$ is
$3H\dot{\phi}+\dfrac{\mathrm{d}V_{\mathrm{eff}}}{\mathrm{d}\phi}\simeq 0\,,$ (21)
which can be integrated by quadrature together with the Friedman equations
$\phi(t)=\phi_{0}+\alpha^{-1}\log\left(1-\dfrac{t}{t_{\mathrm{end}}}\right)\,,$ (22)
$a(t)=e^{H_{0}t}\,\left(1-\dfrac{t}{t_{\mathrm{end}}}\right)^{3/4}\,,$ (23)
$H(t)=H_{0}\,-\dfrac{3}{2}\,t_{\mathrm{end}}^{-1}\,\left(1-\dfrac{t}{t_{\mathrm{end}}}\right)^{-1}\,,$ (24)
where
$t_{\mathrm{end}}=\dfrac{3}{4}\,H_{0}^{-1}\,e^{\alpha\phi_{0}}\,,\quad H_{0}\simeq\dfrac{M_{p}^{4}}{\sqrt{12}(\xi_{s}+\xi_{h}^{2}/\lambda)}\,,$ (25)
is the time that the effective inflation ends and $H_{0}$ is the scale Hubble parameter during inflation, with $\phi_{0}\sim 5.7M_{p}$ is the value of $\phi$ at the beginning of inflation. It is clear from slow-roll solutions that during inflation we have the usual quasi-De Sitter solutions for $\phi$, $a$, and $H$. The effective single-field inflation ends when the slow roll parameter $\epsilon$ is $\epsilon(t_{\mathrm{end}})=1$, which happens at the final values
$\phi(t_{\mathrm{end}})$ $=$ $\alpha^{-1}\log\left(1+\dfrac{2}{\sqrt{3}}\right)\approx 0.92M_{p}\,,$ (26)
$H_{e}$$\approx$ $0.54\,H_{0}\,.$ (27)
3 Scalar fields in the Einstein frame
In this section, we study spectator fields present in the Higgs-$R^{2}$ model such as free scalar fields, initially assuming no direct coupling between the matter fields and the inflationary sector in the Jordan frame. In other words, we get started in the Jordan frame and after conformal transformation, the matter sector is gravitationally coupled to the inflationary sector. In the Jordan frame our model is given by
$S_{J}=\int\mathrm{d}x^{4}\sqrt{-g_{J}}\left[\dfrac{1}{2}(M_{p}^{2}+\xi_{h}H^{2})R_{J}+\dfrac{\xi_{s}}{4}\,R_{J}^{2}-\dfrac{1}{2}g^{\mu\nu}_{J}\,\partial_{\mu}h\,\partial_{\nu}h-\dfrac{\lambda}{4}\,h^{4}\right]+S_{\mathrm{matt}}[\psi_{i}]\,$
where $\psi_{i}$ are the different matter fields that are present during and after inflation. In this work, we will consider a massive spectator scalar field $\chi$ that only interacts gravitationally. The action of such a field is given by
$S_{J}[\chi]=\int\mathrm{d}x^{4}\sqrt{-g_{J}}\left(-\dfrac{1}{2}g^{\mu\nu}_{J}\partial_{\mu}\chi\partial_{\nu}\chi-\dfrac{m_{\chi}^{2}}{2}\chi^{2}\dfrac{1}{2}\,\xi_{\chi}\,R_{J}\,\chi^{2}\right)\,.$ (29)
The dimensionless parameter $\xi_{\chi}$ represents the non-minimal coupling to the gravity sector and may take any value; however, studies based on electroweak vacuum stability [40] suggest that the value of non-minimal coupling constant should not exceed $10$. Moreover there are two interesting special cases: $\xi_{\chi}=0$ and $\xi_{\chi}=1/6$. The value $\xi_{\chi}=0$ is called the case of minimal coupling. while $\xi_{\chi}=1/6$ corresponds to conformal coupling. The first case corresponds to a free scalar field theory and the second one ensures the conformal invariance of the model if, in addition, $m_{\chi}=0$.
Although there is no direct coupling between matter fields and the inflationary sector in the first case, both scalaron $\phi$ and Higgs field $h$ gravitationally interact with every matter field present during inflation and after inflation, and a gravitational particle production process may occur. In the second one, free scalar $\chi$ is coupled to the Ricci scalar and the particle production may also occur as long as the matter fields are not conformally invariant, namely, if the scalar field is massive $m_{\chi}\neq 0$. These kinds of models, in which the scalar fields interact only gravitationally, are especially interesting. They can explain the particle production process, with the gravitational field responsible for creating $\chi$ particles. This process of particle creation due to gravitational interactions has implications for the reheating process [17, 41, 42]. It also plays a role in gravitational dark matter production [43, 44, 45, 12, 11, 9, 46].
What follows is to perform the conformal transformation (4) to the action (29) to change to the Einstein frame. Since $\chi$ is a spectator field, it remains energetically subdominant, $\rho_{\chi}\ll 3H^{2}M_{p}^{2}$, throughout the entire background evolution. Therefore, we can apply the same conformal transformation of the metric (4) as in the previous section. By doing so, we avoid considering back-reaction effects on the space-time geometry and the background dynamics. In the Einstein frame, we obtain the following action
$S_{E}[\tilde{\chi}]=-\dfrac{1}{2}\int\mathrm{d}^{4}x\,\sqrt{-g}\left(g^{\mu\nu}\partial_{\mu}\tilde{\chi}\partial_{\nu}\tilde{\chi}+\xi_{\chi}R\tilde{\chi}^{2}\right.\\[8.5359pt] +\left.e^{-\alpha\phi}m_{\chi}^{2}\tilde{\chi}^{2}\right)+3\alpha\left(\xi_{\chi}-\dfrac{1}{6}\right)\int\mathrm{d}^{4}x\,\sqrt{-g}\left[\tilde{\chi}\,g^{\mu\nu}\partial_{\mu}\tilde{\chi}\partial_{\nu}\phi\right.\\[8.5359pt] \left.+\dfrac{\alpha}{4}\tilde{\chi}^{2}g^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi\right]\,,$ (30)
where $\chi\to\tilde{\chi}=\Psi^{-1}\chi$ is the re-scaled scalar field and the Ricci scalar in the Einstein frame is given by $R=6(\dot{H}+2H^{2})$. We can perform another conformal transformation which involves the scalar field $\tilde{\chi}$ leading to another Einstein frame to remove the non-minimal coupling constant $\xi_{\chi}$. However, we are only interested in the situation where $\chi$ is a spectator field and both scalaron field $\phi$ and Higgs field are responsible for the dynamics during and after inflation. One can see that the conformal transformation (4) induces interaction terms between the scalaron $\phi$, the Higgs field $h$, and the free scalar $\tilde{\chi}$. We can identify the interaction terms from the following action
$S_{\mathrm{int}}=3\alpha\left(\xi_{\chi}-\dfrac{1}{6}\right)\int\mathrm{d}^{4}x\,\sqrt{-g}\left[\tilde{\chi}\,g^{\mu\nu}\partial_{\mu}\tilde{\chi}\partial_{\nu}\phi+\dfrac{\alpha}{4}\tilde{\chi}^{2}g^{\mu\nu}\partial_{\mu}\phi\,\partial_{\nu}\phi\right]\,.$ (31)
This action describes the interaction between the background fields $\phi$ and $h$ with the massive spectator field $\tilde{\chi}$. We can see that even with $\xi_{\chi}=0$ this interaction does not disappear, unless the field $\tilde{\chi}$ is conformal invariant ($m_{\chi}=0\,,\xi_{\chi}=1/6$). In the paper [8], the authors study the gravitational production of dark matter in the $R^{2}$ inflation model and show that the scalaron $\phi$ can decay perturbatively in dark matter particles $\chi$ with masses $m_{\chi}<M/2$, where $M\simeq 1.3\times 10^{-5}M_{p}$ is the mass term of the scalaron. For heavy particles, perturbative particle production is kinematically forbidden. However, the production of scalar particles is not limited to perturbative methods; there may be particle production due to the expansion of the universe and the dynamics of the fields involved during the inflationary epoch. In our case, the classical time-dependent background is defined by the evolution of the scale factor $a(t)$ and the evolution of the scalar fields $\phi$ and $h$. That is, the matter-free scalar field $\chi$ propagates in a time-dependent classical background, whose dynamics are given by the evolution of the $\phi$ and $h$ fields of the Higgs-$R^{2}$ model. To study this process it is necessary to find the equation of motion for the spectator field $\tilde{\chi}$ and track the mode evolution during and after inflation. Varying the action (30) respect to $\tilde{\chi}$ yields the following equation of motion
$(\square-e^{-\alpha\phi}\,m_{\chi}^{2}-\xi_{\chi}\,R)\tilde{\chi}+\left[\dfrac{\alpha}{2}\,\left(1-6\xi_{\chi}\right)\square\phi-\dfrac{\alpha^{2}}{4}\left(1-6\xi_{\chi}\right)(\partial\phi)^{2}\right]\tilde{\chi}=0\,.$ (32)
Notice that the background dynamics is explicit in the equation of the spectator field $\tilde{\chi}$ due to the conformal transformation. This induces a time-dependent effective mass term, $m_{\mathrm{eff}}^{2}(t),$ for the spectator field $\tilde{\chi}$. Consequently, in the Einstein frame, the model presents an interaction term with the dynamics of the classical background through the Ricci scalar, the scale factor $a$, and the dynamics of the fields $\phi$ and $h$. As a result, studying the solutions is generally highly complex from an analytical perspective; hence a numerical treatment of the background dynamics is necessary for a comprehensive analysis of mode evolution, especially after inflation when both fields $\phi$ and $h$ exhibit a strong interacting anharmonic oscillatory behavior, as one can see in the figure 2. By numerically determining the background evolution of the scalaron, the Higgs field, and the scale factor, we can compute the Ricci scalar using $R=6(\dot{H}+2H^{2})$, ensuring that the dynamics are well-determined. However, during effective inflation along one of the potential’s valleys, we can approximate the background dynamics using the single-field equation of motion (20).
4 Mode evolution
In this section, we numerically solve the equation of motion (32) along with the background equations (16), (17), (1) to obtain the mode functions $\chi$ for calculating the number density of particles produced, $n_{k}$. The equation of motion for $\tilde{\chi}$ considering a FLRW universe (1) can be written in terms of the cosmic time $t$ as
$\ddot{\tilde{\chi}}+3H\dot{\tilde{\chi}}+\Omega^{2}(t)\tilde{\chi}=0\,,$ (33)
where $\Omega^{2}(t)$ is a background variables dependent and is given by
$\Omega^{2}(t)=e^{-\alpha\phi}m_{\chi}^{2}-a^{-2}\nabla^{2}+\xi_{\chi}R\\[5.69046pt] +3\alpha\left(\dfrac{1}{6}-\xi_{\chi}\right)(\ddot{\phi}+3H\dot{\phi})-\dfrac{3\alpha^{2}}{2}\left(\dfrac{1}{6}-\xi_{\chi}\right)\dot{\phi}^{2}\,.$ (34)
Rescaling the field $\tilde{\chi}\to a^{-3/2}(t)\,\tilde{\chi}(t,\mathbf{r})$ and using the background equations of motion (16), (19) the equation for the scalar field $\tilde{\chi}$ becomes a oscillator-like equation with time-dependent frequency
$\ddot{X}(t,\mathbf{r})+\omega^{2}(t)\,X(t,\mathbf{r})=0\,,$ (35)
where $\omega^{2}(t)$ is given by
$\omega^{2}(t)=\Omega^{2}(t)-\dfrac{9}{4}H^{2}-\dfrac{3}{2}\dot{H}\,.$ (36)
Particle production in the early universe is a process of quantum field theory in curved spacetime, where the gravitational field is not quantized, but the free scalar field $\chi$ is quantized. Therefore, the production of $\chi$ particles is a quantum effect [3, 4, 5]. We proceed by quantizing the spectator field in the Heisenberg picture so that the $\chi$ field and its conjugate momentum $\pi_{\chi}=\dot{\chi}$ are treated as field operators. In Fourier space, the rescaled field can be written as
$X(t,\mathbf{r})=\int\dfrac{\mathrm{d}^{3}\mathrm{k}}{(2\pi)^{3}}\left(\hat{a}_{\mathbf{k}}\,X_{k}(t)\,e^{-i\mathbf{k}\cdot\mathbf{r}}+\hat{a}^{\dagger}_{\mathbf{k}}\,X^{*}_{k}(t)\,e^{i\mathbf{k}\cdot\mathbf{r}}\right)\,,$
where $\mathbf{k}$ is the comoving wave number and $\hat{a}_{\mathbf{k}}$, $\hat{a}^{\dagger}_{\mathbf{k}}$ are the creation and annihilation operators respectively, which obey the standard commutation relations at equal times $[a_{\mathbf{k}}\,,a^{\dagger}_{\mathbf{k^{\prime}}}]=(2\pi)^{3}\delta^{(3)}(\mathbf{k}-\mathbf{k}^{\prime})$ and $[a_{\mathbf{k}}\,,a_{\mathbf{k^{\prime}}}]=[a^{\dagger}_{\mathbf{k}}\,,a^{\dagger}_{\mathbf{k^{\prime}}}]=0$. The field equation for Fourier mode $X_{k}(t)$ is given by
$\ddot{X}_{k}(t)+\omega_{k}^{2}(t)X_{k}(t)=0\,,$ (38)
with
$\omega_{k}^{2}(t)=\left(\dfrac{k}{a}\right)^{2}+m_{\mathrm{eff}}^{2}(t)\,,$ (39)
and the following effective mass term
$m_{\mathrm{eff}}^{2}(t)=e^{-\alpha\phi}m_{\chi}^{2}+\xi_{\chi}\,R-\dfrac{9}{4}H^{2}-\dfrac{3}{2}\dot{H}+3\alpha\left(\xi_{\chi}-\dfrac{1}{6}\right)(V_{\phi}-\alpha M_{p}^{2}\dot{H})\,.$ (40)
The equation for $X_{k}$ describes an oscillator with time-dependent frequency $\omega_{k}^{2}(t)$, hence to track the evolution of the modes $X_{k}$ during and after inflation, it is necessary to analyze $\omega_{k}^{2}(t)$. One can see that during the quasi De Sitter phase during inflation when the effective single field inflation is valid, the effective mass is simplified because of $\dot{H}\to 0$. In this case, one can see that the effective mass takes the form
$m_{\mathrm{eff}}^{2}(t\ll t_{\mathrm{end}})\simeq e^{-\alpha\phi}m_{\chi}^{2}+12\left(\xi_{\chi}-\dfrac{3}{16}\right)H^{2}+3\alpha\left(\xi_{\chi}-\dfrac{1}{6}\right)V_{\phi}\,.$ (41)
We can see that the mass term is exponentially suppressed during inflation when $\phi\gg M_{p}$, making the mass of the spectator field much smaller than the Hubble parameter at the end of inflation, $m_{\chi}\ll H_{\mathrm{end}}$. However, after inflation, when the scalaron takes values $\phi\ll M_{p}$, the free scalar $\tilde{\chi}$ becomes super heavy, $m_{\chi}\gg H_{\mathrm{end}}$, due to the exponential dependence on the scalaron $\phi$. Then at the beginning of inflation, when $t\ll t_{\mathrm{end}}$ the Ricci scalar is $R\simeq 12H^{2}$ and the dominant terms in (39) are $k^{2}/a^{2}$, $H^{2}$ and $12\xi_{\chi}H^{2}$. However, towards the end of inflation, the term $\dot{H}$ cannot be neglected, as after inflation, $\dot{H}$ and the background oscillations dominate the dynamics. The term $V_{\phi}$ regulates the oscillations of the effective mass after inflation, but it is small compared to $H^{2}$ and $\dot{H}$. The contribution of each term depends on the moment at which it is considered. The mass term benefits from exponential growth when $\phi\ll M_{p}$, as we anticipated earlier. Therefore, after inflation, the dominant terms are the mass term and $\dot{H}$. Additionally, if the coupling constant is $\xi_{\chi}<1/6$, the effective mass could be negative during inflation. In such a case, the frequency $\omega_{k}^{2}(t)$ can become negative for super-horizon modes $k\ll aH$, leading to an exponential mode growth $\tilde{\chi}\sim e^{|\omega_{k}|t}$ during inflation due to tachyonic instability. This regime may result in a very efficient amplification of particle production [41, 33]. However, for large enough modes $k$, namely
$k^{2}>k_{*}^{2}=-m_{\mathrm{eff}}^{2}(t)\,,$ (42)
have positive frequency $\omega_{k}^{2}(t)>0$ even if $m_{\mathrm{eff}}^{2}<0$, which corresponds to sub-horizon regime $k\gg aH$. This occurs when the modes are well inside the horizon at the beginning of inflation. In the figure 3 we show the evolution of $\omega_{k}^{2}(t)$ for a non-minimal coupling $\xi=0$. Before $t_{\mathrm{cross}}$, the frequency is dominated by the decay of the physical mode $k/a\sim e^{-Ht}$. Once the mode crosses the horizon $(k\ll aH)$, $\omega_{k}^{2}(t)$ remains almost constant, and its value is negative since the coefficient of the dominant term $H^{2}$ is negative. For $t\gtrsim t_{\mathrm{end}}$, the frequency increases and oscillates after inflation as the background terms oscillate around the minimum of the potential (preheating).
Figure 3: Time-dependent frequency as a function of cosmic time t for a mass $m_{\chi}=10^{-2}H_{\mathrm{end}}$, a non-minimal coupling $\xi_{\chi}=0$, and $k\gg aH$ for 5 e-folds before horizon crossing. We set $N_{e}=60$ before the end of inflation as the moment when the mode crosses the horizon.
One can also observe growth due to the exponential dependence of the scalaron in the mass term toward the end of inflation. To numerically solve equation (38) along with the background equations, we must specify suitable initial conditions that ensure the modes are within the horizon at early times. In the asymptotic limit $t\to 0$ and according to the behavior of $\omega_{k}^{2}(t)$ at the beginning of inflation, the dominant term is $k/a$. In this limit, the modes are sub-horizon $k\gg aH$, and the modes exhibit Minkowski-like behavior. This motivates choosing the Bunch-Davies vacuum as the initial condition [14]
$\lim_{t\to 0}X_{k}(t)=\dfrac{1}{\sqrt{2\omega_{k}}}\,e^{-i\omega_{k}t}\,.$ (43)
[…]
5 Description of particle production
Once we have established the behavior of mode functions $X_{k}$ during and after inflation for some values of spectator mass, the non-minimal coupling $\xi_{\chi}$ and comoving wavenumber $k$, it turns to analyze the non-perturbative particle production. The production of $\chi$ particles is due to the time dependence of the background variables, in particular, the non-adiabatic behavior of the frequency $\omega_{k}(t)$ that induces a mix among negative- and positive-frequency [3]. The adiabaticity condition is given by the adiabatic parameter $A_{k}$ defined as
$A_{k}(t)=\dfrac{\dot{\omega}_{k}}{\omega_{k}^{2}}\,.$ (44)
The adiabatic evolution of $\omega_{k}(t)$ under the WKB approximation of the modes $X_{k}(t)$ guarantees that the positive-frequency mode at early times evolves into positive-frequency modes at late times. That is, for asymptotically early times $t\to 0$, the modes $X_{k}(t)$ evolve as [47]
$\lim_{t\to 0}X_{k}(t)=\dfrac{e^{-i\int^{t}\omega_{k}(t^{\prime})\mathrm{d}t^{\prime}}}{\sqrt{2\omega_{k}(t)}}\,.$ (45)
This behavior usually happens at early times $t_{i}\to 0$ and at sufficiently late times $t_{f}\to+\infty$, but for intermediate times $t_{i}<t_{*}<t_{f}$ during the early universe, the evolution of the modes is no longer adiabatic $|A_{k}|\gg 1$. During inflation, the positive-frequency solution corresponds to the Bunch-Davies vacuum (43), as described in the previous section. This evolution remains valid as long as the parameter $|A_{k}|\ll 1$ at all times. While this condition is met, no particle production occurs, as adiabatic evolution is ensured, and there is no mixing of negative- and positive-frequency modes. From figure 3, we can see that the evolution of $\omega_{k}(t)$ is smooth until the end of inflation. At that point, the condition $|A_{k}|\ll 1$ is strongly violated, leading to an enhancement in particle production, and gravitational particle production is enhanced at the time $t_{*}$, corresponding to the end of inflation. We show the non-adiabaticity evolution in the figure 4 parametrized by the adiabatic parameter $|A_{k}(t)|$. The mixture of positive and negative frequency modes at time $t_{\mathrm{end}}$ can be represented as the sum of the solutions $\exp(\pm i\int^{t}\omega_{k}(t^{\prime})\mathrm{d}t^{\prime})$ using the WKB approximation [14, 5]
$X_{k}(t)=\dfrac{\alpha(t)}{\sqrt{2\omega_{k}(t)}}\,e^{-i\int^{t}\omega_{k}(t^{\prime})\mathrm{d}t^{\prime}}+\dfrac{\beta(t)}{\sqrt{2\omega_{k}(t)}}\,e^{i\int^{t}\omega_{k}(t^{\prime})\mathrm{d}t^{\prime}}\,,$
where $\alpha_{k}(t)$ and $\beta_{k}(t)$ are the time-dependent Bogolyubov coefficients. Comparing with the Bunch-Davies initial condition (45), one see thats $\alpha_{k}(t_{0})=1$ and $\beta_{k}(t_{0})=0$.
Figure 4: An example of non-adiabatic mode evolution leading the gravitational particle production of $\chi$ particles described by the adiabaticity parameter $|A_{k}(t)|$ for $m_{\chi}=10^{-2}H_{\mathrm{end}}$ and $\xi_{\chi}=1/6$. At the end of inflation, one sees an abrupt increase on $|A_{k}(t)|$ and non-adiabaticity is induced, leading to gravitational particle production $n_{k}(t)\neq 0$. At late times, when the stabilization occurs, we recover the adiabaticity evolution.
In other words, gravitational particle production occurs at very early times $t_{\mathrm{in}}\to 0$ and at late times $t_{f}\to\infty$ if the initial and final vacuum states are not the same. We define the initial vacuum state with the corresponding creation-annihilation operators $\hat{a}^{\dagger}_{\mathbf{k}}$ and $\hat{a}_{\mathbf{k}}$ by
$\hat{a}_{\mathbf{k}}\left|0\right>_{\mathrm{early}}=0\,.$ (47)
The final state will have a different set of creationist-annihilation operators, and thus a different vacuum state. $\hat{a}_{\mathbf{k}}$ it is related to the late-time operators $\hat{b}^{\dagger}_{\mathbf{k}}$ and $\hat{b}_{\mathbf{k}}$ by a Bogolyubov tranformation
$\hat{a}_{\mathbf{k}}=\alpha^{*}_{\mathrm{k}}\hat{b}_{\mathbf{k}}-\beta^{*}_{\mathrm{k}}\hat{b}^{\dagger}_{\mathrm{k}}\,,$ (48)
where $\alpha_{\mathbf{k}}$ and $\beta_{\mathbf{k}}$ are the time-independent Bogolyubov coefficients. The comoving particle number density $n_{\chi}(t)$ is related to the coefficient $\beta_{k}$ of the Bogolyubov transformations which relate the vacuum states at the asymptotic limits $t\to 0$ and $t\to\infty$ and coincides with $\beta_{k}(t)$ at late times. We can compute $\beta_{k}(t)$ by solving (38) and evaluating
$|\beta_{k}(t)|^{2}=\dfrac{\omega_{k}}{2}|X_{k}|^{2}+\dfrac{|\dot{X}_{k}|^{2}}{2\omega_{k}}-\dfrac{1}{2}\,.$ (49)
The physical number density $n_{\chi}(t)$ in the late-time limit $t\to\infty$ is given by
$a^{3}(t)\,n_{\chi}(t)=\int_{0}^{\infty}\mathrm{d}\log k\,n_{k}(t)\,,\quad n_{k}(t)=\dfrac{k^{3}|\beta_{k}(t)|^{2}}{2\pi^{2}}\,,$
where $n_{k}(t)$ is the comoving number density of particles with comoving wavenumber $k$. Notice that the expression (49) is only valid when $\omega_{k}^{2}(t)>0$, so the comoving number density of particles created is valid at times when the frequency (39) is positive.
5.1 Numerical results
What follows is to find the spectrum of $\chi$ particles and their evolution in the cosmic time $t$. Hence we numerically solve (38), constructed the Bogolyubov coefficient $|\beta_{k}(t)|$, and compute the spectrum $n_{k}$ using (50). In the previous section, we described the procedure to numerically evaluate the set of equations to ensure that each mode $k$ is within the horizon at the beginning of the evolution, which allows us to use the Bunch-Davies vacuum (43) as an initial condition of the mode equation for $X_{K}(t)$. The initial conditions for the background are set at time $t_{\mathrm{ini}}$, corresponding to the moment when the modes are within the horizon. For example, for $N_{*}=60$ at the time of crossing, the initial time corresponds to $t_{\mathrm{ini}}\approx 28.712$ in units of $[10^{-5}M_{p}^{-1}]$, and the corresponding mode $k=aH$ will be the value of $aH$ at time $t_{\mathrm{ini}}$. Then, the Bogolyubov coefficient $|\beta_{k}(t)|^{2}$ is calculated using (49), and the comoving number density of gravitational particle production is determined by $n_{k}(t)$.
5.1.1 Conformal coupling $\xi_{\chi}=1/6$
We first present the results for conformally coupled particles ($\xi_{\chi}=1/6$). As an example, we present the time evolution of the comoving number of particles created $n_{k}(t)$ as a function of time for two different cases: light particles $m_{\chi}<H_{\mathrm{end}}$ and heavy particles $m_{\chi}\gtrsim H_{\mathrm{end}}$. We should note that the expression for $\beta_{k}(t)$ is only valid when $\omega_{k}^{2}(t)>0$ , we point out that $n_{k}(t)$ only makes physical sense at late times $t\to\infty$ [5]. The frequency (39) for scalar conformally coupled turns out to be
$\omega_{k}^{2}(t)=\left(\dfrac{k}{a}\right)^{2}+e^{-\alpha\phi}m_{\chi}^{2}+\dfrac{R}{6}-\dfrac{9}{4}H^{2}-\dfrac{3}{2}\dot{H}$ (51)
which can be negative for $k\ll aH$. For this reason, we take into account the sub-horizon evolution of the modes $X_{k}(t)$ and we take $k$ deep inside the horizon at the beginning of the simulation. In the figure, we present two cases: $m_{\chi}=10^{-5}H_{\mathrm{end}}$ and $m_{\chi}=10^{-1}H_{\mathrm{end}}$. In both panels of figure 5, we observe that the number of particles created saturates at the end of inflation, stabilizing and becoming constant at late times $t\to\infty$. This growth in $n_{k}$ is due to the non-adiabaticity of $\omega_{k}$ at the end of inflation, as can be seen in figure 4.
Figure 5: Comoving number density of particles, plotted on a logarithmic scale, as a function of cosmic time $t$ for two different values of the spectator mass $m_{\chi}$, with the same mode $k$, corresponding to $N_{k}=60$. The graph depicts the evolution of the occupation number $n_{k}/k_{\mathrm{end}}^{3}$, where $k_{\mathrm{end}}=a_{\mathrm{end}}H_{\mathrm{end}}$ denotes the mode at the end of inflation
The oscillations of $n_{k}(t)$ after inflation arise from resonances occurring at specific moments between the oscillations of the modes $X_{k}$ and the frequency itself. During this epoch, particle production is driven primarily by the oscillatory behavior of the scalaron and the Higgs field, as illustrated in figure 2, through the term $\dot{H}$ and the Ricci scalar $R$ via the non-minimal coupling $\xi_{\chi}=1/6$. Additionally, we observe that particle production near the end of inflation is most efficient for light particles. In both cases, the particle number is well-defined, grows from the end of inflation, and stabilizes at late times, as expected.
Figure 6: Numerical $n_{k}/k_{\mathrm{end}}^{3}$ vs $k/k_{\mathrm{end}}$ spectra of conformally-coupled scalars for differente light masses $m_{\chi}$. One can see that the UV tail is suppressed for this range of masses.
Figure 7: Right: Comoving number density of particles as a function fo cosmic time $t$. Left: Numerical $n_{k}/k_{\mathrm{end}}^{3}$ vs $k/k_{\mathrm{end}}$ spectra of conformally-coupled scalars for differente large masses $m_{\chi}\gtrsim H_{\mathrm{end}}$. One can see that the UV tail is suppressed for this range of masses.
We have numerically computed the spectra of the comoving number density $n_{k}(t)$ as a function of $k/k_{\mathrm{end}}$ for a range of light masses varying from $10^{-5}H_{\mathrm{end}}$ to $10^{-1}H_{\mathrm{end}}$. Our findings are summarized in figure 6. From the figure, one can observe several remarkable features. Conformally coupled scalars show a spectra that are strongly peaked at some value $k_{\mathrm{peak}}$. For IR modes $(k\ll k_{\mathrm{end}})$, we see that the spectrum is blue-tilted, reaching a peak at some intermediate value $k_{\mathrm{peak}}>k_{\mathrm{end}}$, and then is suppressed for UV modes $(k\gg k_{\mathrm{end}})$ due to inflation, that is, those that remain within the horizon during inflation. IR modes leave the horizon during inflation and re-enter later during the reheating epoch after inflation. Therefore, the dynamics of the $X_{k}(t)$ modes are dictated by the effective mass term (41) with a quasi-De Sitter spacetime as background with $H\simeq\mathrm{constant}$. It is important to note that the UV modes do not cross the horizon during inflation. In place, these modes are mainly produced during the transition to the matter-dominated era and in general during the reheating era. The suppression of UV modes occurs because the $\chi$ scalars become supermassive at the end of inflation due to the exponentially scalaron-dependence mass, causing the effective mass to become substantial. As a result, gravitational particle production becomes less efficient, leading to a significant suppression in the particle density number $n_{k}(t)$. The mass of the field plays a key role in determining the amplitude of $n_{k}$ at the peak momentum $k_{\mathrm{peak}}$, which systematically decreases as $m_{\chi}$ increases [43].
For large masses $m_{\chi}\gtrsim H_{\mathrm{end}}$, the situation is quite similar. As the mass increases, the particle production tends to decrease due to the increase of the effective mass. This can be seen in the left panel of figure 7 for a mass of $m_{\chi}=H_{\mathrm{end}}$ and for a mode crossing the horizon at $N_{k}=60$. In the rigt panel, we present the comoving spectra of particle production for $m_{\chi}\gtrsim H_{\mathrm{end}}$. We find that the behaviour for UV modes is different than in the case of light masses. The location of the peak $k_{\mathrm{peak}}$ is different as a function of the mass value, once $k_{\mathrm{end}}$. While $k_{\mathrm{peak}}>k_{\mathrm{end}}$ as in the case of light particles, the entry to the horizon, and hence the tail suppression of the UV modes is different. This is largely due to the possible resonances that exist for large masses, as the mass term $e^{-\alpha\phi}$ dominates after inflation when $\phi<M_{p}$, which can excite different modes after inflation.
Figure 8: The comoving spectrum of gravitational scalar particle production $n_{k}$ for $\xi_{\chi}=0$ as a function of rescaled horizon modes $k/k_{\mathrm{end}}$. One can see that the spectrum is nearly scale-invariant for masses $m_{\chi}\lesssim H_{\mathrm{end}}$ and blue-tilted behavior for larger masses
5.1.2 Minimal coupling $\xi_{\chi}=0$
For the case of $\xi_{\chi}=0$, we have a different situation. During inflation, the frequency (39) is given by
$\omega_{k}^{2}(t)=\left(\dfrac{k}{a}\right)^{2}+e^{-\alpha\phi}m_{\chi}^{2}-\dfrac{9}{4}H^{2}+\dfrac{7}{6}\dot{H}-\dfrac{\alpha}{2}V_{\phi}\,.$ (52)
In this case, the derivative potential $V_{\phi}$ and the Hubble parameter $H^{2}$ contribute so that the effective mass $m_{\mathrm{end}}$ is negative. Thus, once the modes cross the horizon and for masses $m_{\chi}\ll H_{\mathrm{end}}$, this may lead to $\omega_{k}^{2}(t)<0$, resulting in tachyonic instabilities, as mentioned in the discussion following equation (41) [41]. These instabilities primarily affect the IR modes of the produced particle spectrum, as the modes outside the horizon during inflation experience tachyonic growth. This growth of the modes $X_{\mathbf{k}}(t)$ outside the horizon is the main driver of the enhanced particle production for minimally coupled scalars compared to conformally coupled scalars. The spectrum of the produced particles is also affected, and its behavior depends on the mass $m_{\chi}$. To obtain the gravitational spectra, we use the same technique to solve the equation numerically: we choose the $k$-mode inside the horizon, approximately 5 e-folds before it crosses the horizon and show the numerical result for $\beta_{k}(t)$ at those times when $\omega_{k}^{2}>0$. In the figure 8, we show the spectrum of created particles for minimally coupled masses. We observe that the spectrum is nearly scale-invariant for IR modes and for masses $m_{\chi}\lesssim H_{\mathrm{end}}$ before reaching the peak at $k_{\mathrm{peak}}$. However, as the mass increases, the spectrum tends to be blue-tilted. In the UV region, all modes decay either as a power law or exponentially.
Figure 9: Comoving number density of particles as a function of cosmic time. A similar behavior can be observed in Figure 6: the number of created particles saturates at the end of inflation and stabilizes, becoming constant at late times
A clear difference exists between the cases $\xi_{\chi}=0$ and $\xi_{\chi}=1/6$. In the frequency $\omega_{k}^{2}(t)$, the Ricci scalar does not appear for $\xi_{\chi}=0$, indicating a significant difference in the behavior of $\omega_{k}$ for different coupling cases. This is because the Ricci scalar exhibits an oscillatory behavior after inflation due to the oscillations of the fields $\phi$ and $h$. In other words, for the case $\xi_{\chi}=1/6$, the Ricci scalar plays an important role in describing the evolution of the modes $X_{k}(t)$, which translates into a different spectrum of created particles. From figure 9 we can see that the particle production at minimal coupling is much larger than the conformally-coupled case. This is mainly due to the behavior of the spectator field $X_{k}(t)$ during inflation. As explained above, the tachyonic instabilities in the case of $\xi_{\chi}=0$ cause the modes $X_{k}(t)$ to grow exponentially. This growth of the super-horizon modes is primarily responsible for the enhanced gravitational particle production for minimally coupled scalars.
6 Discussion and remarks
In this study, we investigated the gravitational and non-perturbative production of free scalar particles, $\chi$, in Higgs-$R^{2}$ inflation. We started with a free spectator field, $\chi$, interacting only through the gravitational term $\xi_{\chi}R$ in the Jordan frame. By applying a conformal transformation to transition into the Einstein frame, direct couplings between the $\chi$ field and the scalaron $\phi$ naturally emerged. Within the context of inflationary dynamics in the Einstein frame, we derived the equation of motion for the $\chi$ field, accounting for its non-minimal coupling to the scalar curvature $R$. We numerically solved this system of equations, imposing Bunch-Davies conditions on the $\chi$ field five e-folds before the modes crossed the horizon, to determine the mode functions of $\chi$. From these solutions, we extracted the Bogolyubov coefficient $\beta_{k}(t)$ to compute the comoving particle density and visualize the resulting spectrum. We explored two distinct scenarios for the non-minimal coupling parameter $\xi_{\chi}$, considering different mass ranges. For $\xi_{\chi}=1/6$, we observed efficient particle production at the end of inflation, driven by the term $e^{-\alpha\phi}m_{\chi}^{2}$, which becomes dominant at this stage. The resulting power spectrum exhibited a blue-tilted behavior, indicating that particle production was dominated by modes near the Hubble radius $(aH)^{-1}$ at the end of inflation, while shorter-wavelength modes were suppressed. As the mass increased in the $\xi_{\chi}=1/6$ case, particle production tended to decrease due to the increase in the effective mass. In contrast, for $\xi_{\chi}=0$, the spectrum remained nearly scale-invariant for masses $m_{\chi}\lesssim H_{\mathrm{end}}$, with a blue tilt emerging for larger masses $m_{\chi}\gtrsim H_{\mathrm{end}}$. For light masses and $\xi_{\chi}=0$, we found that the comoving number of particles created was significantly higher compared to the case of conformally coupled particles, primarily due to the tachyonic growth of the modes $X_{k}(t)$.
We did not attempt a detailed exploration of the parameter space of the Higgs-$R^{2}$ inflation model or the parameters of the scalar field $\chi$. Instead, we focused on the Higgs-$R^{2}$ inflation model, whose observables are consistent with Planck data, and considered the possibility of non-perturbative particle production through the direct couplings that arise between the $\chi$ field and the inflationary sector when changing the conformal frame. We found that inflation driven by the Higgs field and the scalaron can produce scalar particles in the cases $\xi_{\chi}=0$ and $\xi_{\chi}=1/6$ for a specific range of masses. This field could potentially contribute to the abundance of dark matter, though its present-day abundance would need to be calculated. There remain several directions for future research, including the production of spin-$1/2$ and spin-$1$ dark matter, the production of gauge bosons $W^{\pm}_{\mu}$, $Z_{\mu}$, and the implications of particle production on primordial non-Gaussianities in the CMB. It would also be interesting to study the isocurvature constraints for light scalar field masses in this model. We leave these intriguing topics for future investigation.
[…]
Acknowledgements
We are very grateful to Dr. Marcos Garcia for his insightful comments and input, which made it possible to write the code for this article. We would also like to acknowledge Dr. Andrew Long for providing an example of numerical code. The work of F.P. has been sponsored by CONACHyT-Mexico through a doctoral scholarship. This work was also partially supported by CONACHyT- Mexico through the support CBF2023-2024-193